OpenStudy (anonymous):

Use synthetic division to determine whether the number k is an upper or lower bound (as specified) for the real zeros of the function f. k = -1; f(x) = 4x3 - 2x2 + 2x + 4; Lower bound?

OpenStudy (anonymous):

did you do the synthetic division?

OpenStudy (anonymous):

@satellite73 no i dont know how to

OpenStudy (anonymous):

it is hard for me to write it here but i can try if you did not do it

OpenStudy (anonymous):

\[ f(x) = 4x^3 - 2x^2 + 2x + 4\] list the coefficients 4 -2 2 4

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

put a \(-1\) on the side 4 -2 2 4 -1 ______________

OpenStudy (anonymous):

bring down the 4 4 -2 2 4 -1 ______________ 4

OpenStudy (anonymous):

then \(-1\times 4=-4\) 4 -2 2 4 -1 -4 ______________ 4

OpenStudy (anonymous):

add and get \(-6\) 4 -2 2 4 -1 -4 ______________ 4 -6

OpenStudy (anonymous):

\(-1\times -6=6\) 4 -2 2 4 -1 -4 6 ______________ 4 -6

OpenStudy (anonymous):

add and get \(8\) 4 -2 2 4 -1 -4 6 ______________ 4 -6 8

OpenStudy (anonymous):

finally \(-1\times 8=-8\) 4 -2 2 4 -1 -4 6 -8 ______________ 4 -6 8

OpenStudy (anonymous):

add and get \(-4\) 4 -2 2 4 -1 -4 6 -8 ______________ 4 -6 8 4

OpenStudy (anonymous):

now look only at the bottom row 4 -6 8 -4

OpenStudy (anonymous):

if the signs were all the same, it would be an upper bound, but they are not however, since the signs alternate + - + - then it IS a lower bound

OpenStudy (anonymous):

not sure how much of that you followed, but that is the method for synthetic division

OpenStudy (anonymous):

thanks so much i have another question if you dont mind helping @satellite73

OpenStudy (anonymous):

go ahead and ask, i will try

OpenStudy (anonymous):

Divide f(x) by d(x), and write a summary statement in the form indicated. f(x) = x4 + 5x3 - 2x2 + 5x - 3; d(x) = x2 + 1

OpenStudy (anonymous):

it is almost impossible to write long division here but i guess we can try it

OpenStudy (anonymous):

x^2 + 1 ( x^4 + 5x^3 -2x^2 + 5x -3

OpenStudy (anonymous):

\(x^2\) goes in to \(x^4\) \(x^2\) times x^2 x^2 + 1 ( x^4 + 5x^3 -2x^2 + 5x -3

OpenStudy (anonymous):

then \(x^2\times (x^2+1)=x^4+x^2\) x^2 x^2 + 1 ( x^4 + 5x^3 -2x^2 + 5x -3 x^4 +x^2

OpenStudy (anonymous):

subtract and get x^2 x^2 + 1 ( x^4 + 5x^3 -2x^2 + 5x -3 x^4 +x^2 ------------------------- 5x^3 - 3x^2

OpenStudy (anonymous):

ick this is a pain \(x^2\) goes in to \(5x^3\) \(5x\) times x^2 + 5x x^2 + 1 ( x^4 + 5x^3 -2x^2 + 5x -3 x^4 +x^2 ------------------------- 5x^3 - 3x^2

OpenStudy (anonymous):

okay then how do you get the remainder

OpenStudy (anonymous):

and \(5x(x^2+1)=5x^3+5x\) x^2 + 5x x^2 + 1 ( x^4 + 5x^3 -2x^2 + 5x -3 x^4 +x^2 ------------------------- 5x^3 - 3x^2 5x^3 + 5x

OpenStudy (anonymous):

gotta divide

OpenStudy (anonymous):

subtract and get x^2 + 5x x^2 + 1 ( x^4 + 5x^3 -2x^2 + 5x -3 x^4 +x^2 ------------------------- 5x^3 - 3x^2 5x^3 + 5x ----------------------- -3x^2 + 0

OpenStudy (anonymous):

f(x) = (x2 + 1)( x2 + 5x - 3) f(x) = (x2 + 1)( x2 + 5x - 3) - 15x - 9 f(x) = (x2 + 1)( x2 - 5x - 3) - 15x - 9 f(x) = (x2 + 1)( x2 - 5x - 3) These are my answer choices how do i get to this

OpenStudy (anonymous):

well you could always multiply out and see which one works

OpenStudy (anonymous):

or just cheat

OpenStudy (anonymous):

how do you cheat

OpenStudy (anonymous):

you see if you divide you get \(5x^2+5x-3\)

OpenStudy (anonymous):

so no remainder